OpenStudy (kainui):

I'm having trouble integrating the following by parts: S(e^(-x)*sin(4x)dx Whenever I try to integrate I end up in this infinite loop situation. How do I get around this?

OpenStudy (kainui):

\[\int\limits_{}^{}e^{-x}\sin(4x)dx\] To be clear, that's my integral.

OpenStudy (anonymous):

Keep integrating by parts, eventually it will oscillate back to: \[\int\limits_{}^{} e ^{-x}\sin(4x) dx\] and you just subtract both sides by that^

OpenStudy (kainui):

Oh, that seems so simple, I guess I kept thinking I'd have to just go down an endless rabbit hole, but once it's on both sides I can subtract it off. Haha thanks.

OpenStudy (anonymous):

Let: u = e^-x du = (-e^-x) dx dv = Sin(4x) dx v = -Cos(4x) / 4 so: \[\int\limits_{}^{} e ^{-x}\sin(4x) dx = -e ^{-x}(1/4)Cos(4x) - \int\limits_{}^{} -e ^{-x} (-Cos(4x)(1/4))dx\] \[\int\limits_{}^{} e ^{-x}\sin(4x) dx = -e ^{-x}Cos(4x) -(1/4)\int\limits_{}^{} e ^{-x}Cos(4x)dx\] Now integrate by parts again: let u = e^-x du = -e^-x dx dv = Cos(4x)dx v = (1/4) Sin(4x) \[\int\limits\limits_{}^{} e ^{-x}\sin(4x) dx = -e ^{-x}Cos(4x) -(1/4)[e ^{-x}(1/4)Sin(4x)- \int\limits_{}^{}-e ^{-x}Sin(4x)(1/4)dx]\] \[\int\limits\limits\limits_{}^{} e ^{-x}\sin(4x) dx = -e ^{-x}Cos(4x) -(1/16)e ^{-x}Sin(4x)+(1/16) \int\limits\limits_{}^{}e ^{-x}Sin(4x)dx\] Do you see the: \[(1/16) \int\limits\limits_{}^{}e ^{-x}Sin(4x)dx\] you see how it shows up on both sides of the equation? lets bring it to one side: \[\int\limits\limits\limits\limits_{}^{} e ^{-x}\sin(4x) dx -(1/16) \int\limits\limits\limits_{}^{}e ^{-x}Sin(4x)dx = -e ^{-x}Cos(4x) -(1/16)e ^{-x}Sin(4x)\] \[(15/16)\int\limits\limits\limits\limits_{}^{} e ^{-x}\sin(4x) dx = -e ^{-x}Cos(4x) -(1/16)e ^{-x}Sin(4x)\] now just bring the (15/16) back to the other side: \[\int\limits\limits\limits\limits_{}^{} e ^{-x}\sin(4x) dx = (16/15) [-e ^{-x}Cos(4x) -(1/16)e ^{-x}Sin(4x)]\] tada!

OpenStudy (anonymous):

lol guessed you figured it out before i finished typing. Good job!